in carbon monoxide molecules, the carbon and the oxygen atoms are separated by distance 1.2 armstrong. the distance of the centre of mass, from the carbon atom is

Grade:12

2 Answers

Arun
6 years ago
Dear Srujana
 
mass of the C atom = 12 units
mass of the O atom = 16 units
 position of the C atom is taken to be origin.
hence,
position of centre of mass : = 1.2* 16 /(12+16)
 = 12*16/28 = 0.69 Angstrom
 
Regards
Arun (AskIITians Forum Expert)
If you find the solution correct then please approve it & in case of any confusion further please do feel free to write me here. I would be more than happy to help further
Kushagra Madhukar
3 years ago
Dear Student,
Please find the solution to your problem.
 
mass of the C atom, m1 = 12 units
mass of the O atom, m2 = 16 units
 position of the C atom is taken to be origin.
hence,
position of centre of mass, xcm  = (m1x1 + m2x2)/(m1 + m2)
                                                = (12*0 + 1.2* 16) /(12+16)
                                                = 1.2*16/28
                                                = 0.69 Angstrom
 
Thanks and regards,
Kushagra